OpenStudy (anonymous):

find the equation of the tangent line to the graph f(x)=x^3 at the given x= -2

OpenStudy (anonymous):

when x=-2 y=f(x)=(-2)^3=-8 so point is (-2,-8 ) next first find the slope sdlope is equal to derivative differentiating \[\Large m=\frac {dy}{dx}=3x^2\] put x-2 so \[\Large m=3(-2)^2=12\] use point slope form \[\Large y-y_{0}=m(x-x_{0})\] where m=12 x0=-2 and y0=-8 put in the above equation to get the required equation of tangent line .

OpenStudy (anonymous):

I got y=12x+16 as the answer.

OpenStudy (anonymous):

this is correct .